# Neural reproducing kernel Banach spaces and representer theorems for deep networks

### v1

@@ Abstract.
Studying the function spaces defined by neural networks helps to understand the corresponding learning models and their inductive bias. While in some limits neural networks correspond to function spaces that are reproducing kernel Hilbert spaces, these regimes do not capture the properties of the networks used in practice. In contrast, in this paper we show that deep neural networks define suitable reproducing kernel Banach spaces. These spaces are equipped with norms that enforce a form of sparsity, enabling them to adapt to potential latent structures within the input data and their representations. In particular, leveraging the theory of reproducing kernel Banach spaces, combined with variational results, we derive representer theorems that justify the finite architectures commonly employed in applications. Our study extends analogous results for shallow networks and can be seen as a step towards considering more practically plausible neural architectures.
@@ 1. Introduction
Neural networks define functions by composing linear and nonlinear maps in a multi-layer (deep) architecture. While easy to implement, the corresponding models are hard to analyze since they are nonlinearly parameterized. Understanding the properties of the function spaces defined by different neural network architectures can give insights in the corresponding learning models. Further, it can provide indications on the underlying inductive bias, namely what functions can be approximated and learned efficiently by a given class of networks.
@@
In some overparametrized regimes, neural networks can be seen to define Hilbert spaces of functions and in particular reproducing kernel Hilbert spaces (RKHS) [1]. For instance, a classical observation is that shallow networks with infinitely many random units correspond to RKHS, with reproducing kernels depending on the considered nonlinearity [15]. This regime, also known as the Gaussian Process (GP) limit, has connections with models such as random features [19].The limits of more complex, possibly non shallow architectures can also be derived and characterized in terms of RKHS, see e.g. [12]. Another infinite-width limit in which neural networks are described by RKHS is the so-called lazy training regime [9]. In this limit, the network weights evolve little during the optimization and can be well approximated by a linear approximation around a random initialization. In this case as well, the corresponding function spaces are RKHS, and the associated kernel is called neural tangent kernel (NTK) [13]. Again, neural tangent kernels and corresponding RKHS can be derived for a variety of architectures, see e.g. [5].
@@
In fact, neither the above GP/random features limit nor the NTK/lazy training regime seem to capture key aspects of neural networks models [11, 4]. Results for shallow networks suggest that neural networks might favor functions with small norms that are not Hilbertian but rather associated to Banach spaces [2]. In turn, representer theorems associated to such norms allow to characterize finite-width networks commonly used in practice [22, 16, 24, 17, 18]. These observations have sparked considerable interest in understanding the Banach spaces associated to neural networks. Notably, one possibility is to consider extensions of classical splines [28, 27, 26, 25]. Another possibility is to consider reproducing kernel Banach spaces (RKBS) [30, 14], see e.g. [3] and references therein.
In this paper, we develop the latter approach tackling the extension from shallow to deep networks. The study of Banach spaces associated to deep architectures and corresponding representer theorems was started in [18], where deep architectures with ReLU activations and finite rank constraints at each layer are considered. The latter requirement is not natural and is mainly due to technical reasons. Indeed, the finite rank constraint allows for the construction of layers as concatenation of real valued functions studied for shallow networks.
@@
In our study, we propose an approach which allows to consider more general activations and especially to avoid the finite rank constraints. This requires more substantial developments using vector Radon measures to address the challenges posed by potentially infinite-dimensional hidden layers. Our first contribution is to define a reproducing kernel Banach space which describes an infinite-width limit of a deep neural network with an associated norm promoting sparsity. We dub such a space a neural RKBS. Then, we provide a representer theorem for a large class of nonlinearities that shows how optimal neural networks minimizing empirical objectives can be taken to have a finite structure at every layer. This result extends analogous results for shallow networks. It implies that commonly used networks are optimal in the sense that they are solutions of a suitable variational problem.
@@
The rest of the paper is organized as follows. In Section 2 we review the main technical ingredients of our construction, namely reproducing kernel Banach spaces and vector Radon measures. In Section 3 we introduce several notions of RKBS to model functional properties of neural networks, leading to the construction of deep and neural RKBS. In Section 4 we prove our represent theorems on deep and neural RKBS. In Appendix A we collect variational results and extreme point characterizations used to prove our representer theorems.
@@
$B(X,Y)$ bounded linear maps $X\to Y$ $\mathcal{M}(\Theta)$ $\mathcal{M}(\Theta,\mathbb{R})$
$X^{\prime}$ continuous dual of $X$ $\|\mu\|_{\operatorname{TV}}$ total variation norm of measure $\mu$
,${}_{X}\langle\cdot,\cdot\rangle_{X^{\prime}}$ pairing on $X,X^{\prime}$ $\delta_{\theta}$ Dirac delta at $\theta$
$\langle\cdot,\cdot\rangle_{X}$ inner product on $X$ $\mathcal{C}_{0}(\Theta,Y)$ continuous functions $\Theta\to Y$ vanishing at $\infty$
$\|\cdot\|_{X}$ norm on $X$ $\mathcal{C}_{0}(\Theta)$ $\mathcal{C}_{0}(\Theta,\mathbb{R})$
$\mathcal{B}(\Theta)$ Borel $\sigma$-algebra on $\Theta$ $B_{X}(r)$ ball on $X$ of radius $r$
$\mathcal{M}(\Theta,Y)$ vector measures on $\Theta$ with values in $Y$ $\operatorname{Ext}(Q)$ extremal points of $Q$
2. Preliminaries
In this section, we review the notion of an architecture and building blocks of deep neural networks. Our perspective is to cast the construction of neural networks in the theory of reproducing kernel Banach spaces, as functions parametrized by vector measures.
@@ 2. Preliminaries
We start by setting up some notation, considering to the case of fully connected, feed-forward architectures.
@@
Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a (nonlinear) function, $L\geq 1$ an integer and $d=d_{0},d_{1},\dots,d_{L},d_{L+1}=p\geq 1$ a family of integers. A (fully connected feed-forward) neural network from $\mathbb{R}^{d}$ into $\mathbb{R}^{p}$ with activation function $\sigma$, depth $L$ and widths $d_{1},\dots,d_{L}$ is any function $f:\mathbb{R}^{d}\to\mathbb{R}^{p}$ of the form $f(x)=x^{(L+1)}$ where
$\begin{cases}x^{(1)}=W^{(1)}x+b^{(1)}&\in\mathbb{R}^{d_{1}}\\[5.0pt] x^{(\ell+1)}=W^{(\ell+1)}\sigma(x^{(\ell)})+b^{(\ell+1)}&\in\mathbb{R}^{d_{\ell+1}}\qquad\ell=1,\dots,L\end{cases}$ (1)
@@
for some weights $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ and biases (or offsets) $b^{(\ell)}\in\mathbb{R}^{d_{\ell}}$, and where the function $\sigma$ is assumed to apply on vectors component by component. We call a neural network shallow (or a one-hidden layer network) if $L=1$, deep if $L>1$. A neuron is a function of the form $\phi(z)=\sigma(\langle z,w\rangle+b)$. The vector $\sigma(x^{(\ell)})$ for $\ell=1,\dots,L$ is the $\ell$-th hidden layer of the network, hence $d_{\ell}$ is the number of neurons at the $\ell$-th hidden layer.
The input and output dimensions $d,p$ are prescribed by the problem. Given the problem, the user chooses an activation $\sigma$, a depth $L$ and widths $d_{1},\dots,d_{L}$. After this choice, the architecture is fixed, and the parameters $W^{(\ell)},b^{(\ell)}$ are obtained optimizing an empirical loss on the available data. Optimization is thus performed on the set of functions (1), with $(W^{(\ell)},b^{(\ell)})$ in $\mathbb{R}^{d_{\ell}\times d_{\ell-1}}\times\mathbb{R}^{d_{\ell}}$. The set of neural networks with a fixed architecture is parametric, albeit hence nonlinear and high-dimensional. Next, we will define structured nonparametric linear spaces on which optimization has a solution in the set of neural networks (1). This suggests that neural networks may be seen as parametric solutions of a nonparametric model, where the architecture itself can be obtained by training, rather than arbitrarily chosen by the practitioner. To define such a model, we need the notion of a reproducing kernel Banach space and vector measures.
Reproducing kernel Banach spaces
@@
We will assume the following minimal definition, that readily generalizes reproducing kernel Hilbert spaces to a Banach setting. We refer to [30, 14] for a comprehensive overview.
Definition 2.2 (Vector valued reproducing kernel Banach space).
@@
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space. A reproducing kernel Banach space (RKBS) $\mathcal{H}$ on $\mathcal{X}$ with values in $\mathcal{Y}$ is a Banach space of functions $f:\mathcal{X}\to\mathcal{Y}$ with continuous pointwise evaluation, that is, for all $x\in\mathcal{X}$ there is $C_{x}>0$ such that $\|f(x)\|_{\mathcal{Y}}\leq C_{x}\|f\|_{\mathcal{H}}$ for all $f\in\mathcal{H}$. In other words, the map $f\mapsto f(x)$ is in $B(\mathcal{H},\mathcal{Y})$ for all $x\in\mathcal{X}$.
The following proposition characterizes RKBS as Banach feature spaces, that is, spaces parametrized on Banach spaces.
@@
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space. Consider the following statements.
@@
    A space $\mathcal{H}$ of functions $f:\mathcal{X}\to\mathcal{Y}$ is a RKBS.
@@
      =$\mathcal{H}=\{f_{\mu}:\mu\in\mathcal{F}\}$ where $f_{\mu}=\phi(\cdot)\mu$ ;
@@
      $\|f\|_{\mathcal{H}}=\inf\{\|\mu\|_{\mathcal{F}}:\mu\in\mathcal{F},f=f_{\mu}\}$
@@
      =$\mathcal{H}=\{f_{\mu}:\mu\in\mathcal{F}\}$ where ${}_{\mathcal{Y}}\langle f_{\mu}(\cdot),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}={}_{\mathcal{F}}\langle\mu,\psi(\cdot)y^{\prime}\rangle_{\mathcal{F}^{\prime}}$.
@@
      $\|f\|_{\mathcal{H}}=\inf\{\|\mu\|_{\mathcal{F}}:\mu\in\mathcal{F},f=f_{\mu}\}$
Then (a) and (b) are equivalent and each one implies (c). Moreover, if $Y$ is reflexive (in particular Hilbert), then (a), (b) and (c) are all equivalent.
@@
$\phi:\mathcal{X}\to B(\mathcal{F},\mathcal{Y})\qquad\phi(x)f=f(x)\ .$
Then (b)(i) and (b)(ii) of item (b) are clear. Let us prove that (b) implies (a). Clearly $\mathcal{H}$ is a linear space and $\|\cdot\|_{\mathcal{H}}$ is a norm. We then show that the normed space $\mathcal{H}$ is complete. The linear map $\mu\mapsto f_{\mu}$ has kernel $\mathcal{N}=\bigcap_{x\in\mathcal{X}}\ker\phi(x)$. Since $\phi(x)$ is bounded for all $x$, $\ker\phi(x)$ is closed, hence so is $\mathcal{N}$. Thus, $\mathcal{F}/\mathcal{N}$ is a Banach space [20, Theorem 1.41], by construction isomorphic to $\mathcal{H}$, which is therefore complete. Next, we show that point evaluations in $\mathcal{H}$ are continuous. For $f\in\mathcal{H}$, let $\mu\in\mathcal{F}$ such that $f=f_{\mu}$. Then
@@
The implication from (b) to (c) follows easily considering $\psi(x)$ as the dual map of $\phi(x)$. Finally, note that (c)(i) in (c) defines $f_{\mu}$ as a function from $\mathcal{X}$ to $\mathcal{Y}^{\prime\prime}$. Hence, if $\mathcal{Y}$ is reflexive, it defines $f_{\mu}:\mathcal{X}\to\mathcal{Y}$. From here, following the proof of the implication from (b) to (a), one can prove that (c) implies (a). ∎
@@
In view of 2.3, RKBS can be constructed choosing a feature space parametrizing functions. Thinking of a neural network layer as an atomic integration, we will define RKBS parametrized by measures. As hidden layers have vectorial outputs, we need the concept of vector measure [10].
Definition 2.4 (Vector measures).
Let $\Theta$ be a (Hausdorff) locally compact, second countable topological space, and let $\mathcal{Y}$ be a Banach space. A vector (Radon) measure on $\Theta$ with values in $\mathcal{Y}$ is a map $\mu\colon\mathcal{B}(\Theta)\to\mathcal{Y}$ such that, for all $A\in\mathcal{B}(\Theta)$ and all $\{A_{i}\}$ partitions of $A$, [^1] A partition of a Borel set $A$ is a numerable family of Borel sets $\{A_{i}\}$ such that $A_{i}\cap A_{j}=\emptyset$ for all $i\neq j$ and $\bigcup_{i}A_{i}=A$.

### v2

@@ Abstract.
Characterizing the function spaces defined by neural networks helps understanding the corresponding learning models and their inductive bias. While in some limits neural networks correspond to function spaces that are Hilbert spaces, these regimes do not capture the properties of the networks used in practice. Indeed, several results have shown that shallow networks can be better characterized in terms of suitable Banach spaces. However, analogous results for deep networks are limited. In this paper we show that deep neural networks define suitable reproducing kernel Banach spaces. These spaces are equipped with norms that enforce a form of sparsity, enabling them to adapt to potential latent structures within the input data and their representations. In particular, by leveraging the theory of reproducing kernel Banach spaces, combined with variational results, we derive representer theorems that justify the finite architectures commonly employed in applications. Our study extends analogous results for shallow networks and represents a step towards understanding the function spaces induced by neural architectures used in practice.
@@ 1. Introduction
Neural networks define functions by composing linear and nonlinear maps in a multi-layer (deep) architecture. While easy to implement, the corresponding models are hard to analyze since they are nonlinearly parameterized. The study of the function spaces defined by different neural network architectures provides insights into the corresponding learning models. In particular, it provides indications on the underlying inductive bias, namely, which functions can be approximated and learned efficiently by a given class of networks.
@@
In some over-parameterized regimes, neural networks can be seen to define Hilbert spaces of functions and in particular reproducing kernel Hilbert spaces (RKHS) [1]. It is a classical observation that shallow networks with infinitely many random units correspond to RKHS, with reproducing kernels depending on the considered activation function [19]. This regime, also known as the Gaussian Process (GP) limit, has connections with models such as random features [23]. The limits of more complex, possibly non-shallow architectures can also be derived and characterized in terms of RKHS, see e.g. [16]. Another infinite-width limit in which neural networks are described by RKHS is the so-called lazy training regime [10]. In this limit, the network weights evolve little during the optimization and can be well approximated by a linear approximation around a random initialization. Also in this case, the corresponding function spaces are RKHS, and the associated kernel is called neural tangent kernel (NTK) [17]. Again, neural tangent kernels and corresponding RKHS can be derived for a variety of architectures, see e.g. [5].
@@
However, neither the above GP/kernel limit nor the NTK/lazy training regime appear to capture key aspects of neural network models [15, 4] used in practice. Results for shallow networks suggest that neural networks might favor functions with small norms that are not Hilbertian but rather associated with Banach spaces [2]. In turn, representer theorems associated to such norms allow to derive finite-width networks commonly used in practice from variational principles [26, 20, 28, 21, 22]. These observations have sparked interest in understanding Banach spaces associated to neural networks. One possibility is to consider extensions of classical splines [32, 31, 30, 29]. Another possibility is to consider reproducing kernel Banach spaces (RKBS) [34, 18], see e.g. [3] and references therein.
@@
In this paper, we develop the latter perspective tackling the extension from shallow to deep networks. The study of Banach spaces associated to deep architectures and corresponding representer theorems was started in [22], where deep architectures with ReLU activations and finite rank constraints at each layer are considered. The latter requirement is not natural and is mainly due to technical reasons. Indeed, finite rank constraints allow for the construction of layers as concatenation of vector valued functions studied for shallow networks. In our study, we propose an approach that avoids finite rank constraints and allows to consider more general activations. This requires more substantial developments employing vector measures of finite variation to address the challenges posed by potentially infinite-dimensional hidden layers. Our first contribution is to define a reproducing kernel Banach space which describes an infinite-width limit of a deep neural network with an associated norm promoting sparsity. We call such a space a neural RKBS. Then, we provide a representer theorem for a large class of nonlinearities that shows how neural networks minimizing empirical objectives can be taken to have a finite width at every layer. This result extends analogous results for shallow networks. It implies that commonly used networks are optimal in the sense that they are solutions of a suitable variational problem.
@@
The rest of the paper is organized as follows. In Section˜2 we provide background for the study of the paper. In Section˜3 we review the main technical ingredients of our construction, namely reproducing kernel Banach spaces and vector Radon measures. In Section˜4 we introduce deep integral RKBS to model functional properties of deep neural networks, leading to the construction of neural RKBS. In Section˜5 we prove our represent theorems for deep neural networks. In Section˜6, we construct a particular instance of neural RKBS with a countably infinite number of neurons per hidden layer, which permits a more explicit form of the relative representer theorem. In Appendix˜A we collect variational results and extreme point characterizations used to prove our representer theorems. Table 1 summarizes the main notation we use in the paper.
@@
$\mathcal{X},\mathcal{Y}$ Banach spaces $\Theta$ locally compact second contable space
$B(\mathcal{X},\mathcal{Y})$ bounded linear maps $\mathcal{X}\to\mathcal{Y}$ $\mathcal{B}(\Theta)$ Borel $\sigma$-algebra on $\Theta$
$\mathcal{X}^{\prime}$ continuous dual of $\mathcal{X}$ $\mathcal{C}_{0}(\Theta,\mathcal{Y})$ continuous functions $\Theta\to\mathcal{Y}$ vanishing at $\infty$
$_{\mathcal{X}}\langle\cdot,\cdot\rangle_{\mathcal{X}^{\prime}}$ pairing on $\mathcal{X},\mathcal{X}^{\prime}$ $\mathcal{C}_{0}(\Theta)$ $\mathcal{C}_{0}(\Theta,\mathbb{R})$
$\langle\cdot,\cdot\rangle_{\mathcal{X}}$ inner product on the Hilbert space $\mathcal{X}$ $\mathcal{M}(\Theta,\mathcal{Y})$ vector measures on $\Theta$ with values in $\mathcal{Y}$
$\|\cdot\|_{\mathcal{X}}$ norm on $\mathcal{X}$ $\mathcal{M}(\Theta)$ $\mathcal{M}(\Theta,\mathbb{R})$
$B_{\mathcal{X}}(r)$ ball on $\mathcal{X}$ of radius $r$ $\delta_{\theta}$ Dirac delta at $\theta$
$\operatorname{Ext}(Q)$ extremal points of $Q\subset\mathcal{X}$ $\|\cdot\|_{\operatorname{TV}}$ total variation norm
2. Preliminary Discussion
In this section, we provide the necessary background by introducing the notation for deep neural networks, reviewing the infinite-width limit of shallow networks, and presenting a high-level overview of the proposed infinite-width limit for deep architectures.
@@ 2. Preliminary Discussion
We start by setting up some notation and introduce fully connected feed-forward neural networks.
@@
Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a (nonlinear) function, $L\geq 1$ an integer and
$d=d_{0},d_{1},\dots,d_{L},d_{L+1}=p\geq 1$
@@
a family of $L+2$ integers. A function $f:\mathbb{R}^{d}\to\mathbb{R}^{p}$ is called a fully connected feed-forward neural network from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ with activation function $\sigma$, depth $L$ and widths $d_{1},\dots,d_{L}$ if
$f(x)=x^{(L+1)}$
where, for each $x\in\mathbb{R}^{d}$, the vector $x^{(L+1)}\in\mathbb{R}^{p}$ is defined by the following recursive equation
$\begin{cases}x^{(1)}=W^{(1)}x+b^{(1)}&\in\mathbb{R}^{d_{1}}\\[5.0pt] x^{(\ell+1)}=W^{(\ell+1)}\sigma(x^{(\ell)})+b^{(\ell+1)}&\in\mathbb{R}^{d_{\ell+1}},\qquad\ell=1,\dots,L,\end{cases}$ (1)
@@
for some weights $W^{(\ell)}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ and biases (or offsets) $b^{(\ell)}\in\mathbb{R}^{d_{\ell}}$. We call a neural network shallow (or a one-hidden layer network) if $L=1$, deep if $L>1$.
In (1) the activation function $\sigma$ is assumed to be applied on vectors component-wise, and the vector $\sigma(x^{(\ell)})$ for $\ell=1,\dots,L$ is the $\ell$-th hidden layer of the network. The input and output dimensions $d=d_{0},p=d_{L+1}$ are fixed by the problem, and $d_{\ell}$ is the number of neurons (or units) at the $\ell$-th hidden layer. A neuron is a function of the form $\phi(z)=\sigma(\langle z,w\rangle+b)$. Hence, the definition of a network requires specifying an activation function $\sigma$, a depth $L$, and the widths $d_{1},\dots,d_{L}$. These parameters define the network architecture. Given an architecture, the parameters $W^{(\ell)},b^{(\ell)}$ in $\mathbb{R}^{d_{\ell}\times d_{\ell-1}}\times\mathbb{R}^{d_{\ell}}$ are found optimizing an empirical objective function.
The neural networks with a fixed architecture form a subset $\mathcal{F}_{NN}$ of functions from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ parameterized nonlinearly by weights and biases. This is already clear considering scalar-valued shallow neural networks of width $K\in\mathbb{N}$, that is
$\mathcal{F}_{NN}=\left\{f:\mathbb{R}^{d}\to\mathbb{R}:f(x)=\sum_{k=1}^{K}v_{k}\sigma(w_{k}^{\top}x+b_{k})+b,w_{k}\in\mathbb{R}^{d},v_{k},b,b_{k}\in\mathbb{R}\right\}.$ (2)
The nonlinear dependence on the parameters is a key challenge in defining and characterizing the corresponding function spaces. For shallow networks, considering the so called infinite-width limit provides an approach to tackle this question.
Infinite-width limit of shallow neural networks.
In the context of shallow networks, the infinite-width limit corresponds to consider functions parameterized by measures rather than weights, that is
$f_{\mu}(x)=\int_{\mathbb{R}^{d}\times\mathbb{R}}\sigma(w^{\top}x+b)\,{\rm d}\mu(w,b),\qquad\mu\in\mathcal{M}(\mathbb{R}^{d}\times\mathbb{R}).$ (3)
Generalizing further, one can consider an arbitrary input space $\mathcal{X}$ and functions $f$ from $\mathcal{X}$ to $\mathbb{R}$ of the form
$f_{\mu}(x)=\int_{\Theta}\rho(x,\theta)\,{\rm d}\mu(\theta),\qquad\mu\in\mathcal{M}(\Theta).$ (4)
for a suitable locally compact second countable parameter space $\Theta$ and basis function $\rho:\mathcal{X}\times\Theta\to\mathbb{R}$.
We add two observations. First, the expression in (3) needs some care. Indeed, commonly used activation functions are not integrable, and typical parameter spaces $\Theta$ are non-compact, so that the integrals in (3) (and (4)) may not converge. For example, in [3] the activation is multiplied by a smoothing function to ensure integrability. The corresponding function space is shown to be a reproducing kernel Banach space, with norm induced by the total variation norm on measures. Second, the finite shallow networks in the set (2) can be recovered from the infinite-width model (3) by taking finite linear combinations of Dirac deltas as measures. Interestingly, finite networks can be shown to emerge from variational principles. Specifically, if the minimization of an empirical objective is performed over measures, instead of weights, then the optimal solutions are atomic measures. In other words, finite networks are optimal solutions of empirical minimization problems over possibly infinite dimensional networks. We refer for example to [29] for an account of recent works addressing the above topics. As discussed next, our main motivation is studying similar questions in the context of deep networks.
Infinite-width limit of deep neural networks
Previous work on Banach spaces associated with deep architectures, and corresponding representation theorems, focus on layers which are concatenation of of infinite-width shallow neural networks with finite-dimensional outputs. As a result, this approach inherently restricts the deep architecture to finite-rank constraints at every layer [22]. In contrast, our work presents an alternative framework that avoids such constraints and accommodates a broader class of activation functions. This generalization requires addressing the challenges posed by potentially infinite-dimensional hidden layers. To this end, we construct deep architectures as compositions of RKBS of integrable functions valued in infinite-dimensional spaces. More precisely, we begin by considering the direct sum of integral RKBS, forming a linear space, and use this structure to parameterize spaces of composed functions, thereby yielding a corresponding nonlinear function space, which we call a deep integral RKBS. By further specializing this construction to integral functions defined through activation functions, we arrive at what we call a neural RKBS. This formulation provides a natural and rigorous connection with commonly used deep neural network models. Finally, we use these function spaces to derive novel representer theorems for deep networks. These representer theorems characterize the minimization of empirical objective functions over deep RKBS. In particular, they show that finite networks are optimal from a variational perspective. Next, we develop these ideas in detail.
3. Vector measures, integral RKBS and Neural Networks
@@
We introduce the following definition, that readily generalizes vector-valued reproducing kernel Hilbert spaces [8] to a Banach setting. We refer to [34, 18] for an overview.
Definition 3.1 (Vector-valued RKBS).
@@
Let $\mathcal{X}$ be a set and $\mathcal{Y}$ a Banach space. A reproducing kernel Banach space (RKBS) $\mathcal{H}$ on $\mathcal{X}$ with values in $\mathcal{Y}$ is a Banach space such that
  1. (i)
    the elements of $\mathcal{H}$ are functions $f:\mathcal{X}\to\mathcal{Y}$;
  2. (ii)
    the sum and the multiplication by scalars in $\mathcal{H}$ are defined pointwise;
  3. (iii)
    for all $x\in\mathcal{X}$ there is $C_{x}>0$ such that $\|f(x)\|_{\mathcal{Y}}\leq C_{x}\|f\|_{\mathcal{H}}$ for all $f\in\mathcal{H}$.
The first two conditions are equivalent to say that $\mathcal{H}$ is a subspace of $\mathcal{Y}^{\mathcal{X}}$, the vector space of all functions from $\mathcal{X}$ to $\mathcal{Y}$, while the third condition requires that for all $x\in\mathcal{X}$ the pointwise evaluation $f\mapsto f(x)$ is in $B(\mathcal{H},\mathcal{Y})$, the space of bounded linear operators between $\mathcal{H}$ and $\mathcal{Y}$.
Reproducing kernel Hilbert spaces can be characterized in terms of so-called feature maps. In the following proposition we provide an analogous result for RKBS. In this case, feature spaces are Banach spaces.
@@
Let $\mathcal{X}$ be a set, $\mathcal{Y}$ a Banach space and $\mathcal{H}$ a set of functions $f:\mathcal{X}\to\mathcal{Y}$. Consider the following statements.
@@
    The space $\mathcal{H}$ is a RKBS.
@@
      $\mathcal{H}=\{f_{\mu}:\mu\in\mathcal{F}\}$ where $f_{\mu}=\phi(\cdot)\mu$ ;
@@
      $\|f\|_{\mathcal{H}}=\inf\{\|\mu\|_{\mathcal{F}}:\mu\in\mathcal{F},f=f_{\mu}\}$ .
@@
      $\mathcal{H}=\{f_{\mu}:\mu\in\mathcal{F}\}$ where $_{\mathcal{Y}^{\prime\prime}}\langle f_{\mu}(\cdot),y^{\prime}\rangle_{\mathcal{Y}^{\prime}}={}_{\mathcal{F}}\langle\mu,\psi(\cdot)y^{\prime}\rangle_{\mathcal{F}^{\prime}}$ for all $y^{\prime}\in\mathcal{Y}^{\prime}$;
@@
      $\|f\|_{\mathcal{H}}=\inf\{\|\mu\|_{\mathcal{F}}:\mu\in\mathcal{F},f=f_{\mu}\}$ .
Then (a) and (b) are equivalent and each one implies (c). Moreover, if $\mathcal{Y}$ is reflexive (in particular Hilbert), then (a), (b) and (c) are all equivalent.
@@
$\phi:\mathcal{X}\to B(\mathcal{F},\mathcal{Y}),\qquad\phi(x)f=f(x).$
Then (b)(i) and (b)(ii) of item (b) are clear. Let us prove that (b) implies (a). Clearly $\mathcal{H}$ is a linear space and $\|\cdot\|_{\mathcal{H}}$ is a norm. We then show that the normed space $\mathcal{H}$ is complete. The linear map $\mu\mapsto f_{\mu}$ has kernel $\mathcal{N}=\bigcap_{x\in\mathcal{X}}\ker\phi(x)$. Since $\phi(x)$ is bounded for all $x\in\mathcal{X}$, $\ker\phi(x)$ is closed, hence so is $\mathcal{N}$. Thus, $\mathcal{F}/\mathcal{N}$ is a Banach space [24, Theorem 1.41] isomorphic to $\mathcal{H}$ by construction, which is therefore complete. Next, we show that point evaluations in $\mathcal{H}$ are continuous. For $f\in\mathcal{H}$, let $\mu\in\mathcal{F}$ such that $f=f_{\mu}$. Then
@@
The implication from (b) to (c) follows easily considering $\psi(x)=\phi(x)^{t}\in\mathcal{B}(\mathcal{Y}^{\prime},\mathcal{F}^{\prime})$ be the transpose map of $\phi(x)$. Finally, note that (c)(i) in (c) defines $f_{\mu}$ as a function from $\mathcal{X}$ to $\mathcal{Y}^{\prime\prime}$. Hence, if $\mathcal{Y}$ is reflexive, it defines $f_{\mu}:\mathcal{X}\to\mathcal{Y}$. From here, following the proof of the implication from (b) to (a), one can prove that (c) implies (a). ∎
@@
In view of ˜3.2, a RKBS can be constructed choosing a suitable feature space. Thinking of a neural network layer as an atomic integration, we will define RKBS parameterized by measures. Since layers have vectorial outputs, we need the notion of vector valued measure [11].
Let $\Theta$ be a Hausdorff, locally compact, second countable topological space, and let $\mathcal{Y}$ be a Banach space. Recall that a (numerable) partition of a Borel set $A$ is a numerable family of Borel sets $\{A_{i}\}$ such that $A_{i}\cap A_{j}=\emptyset$ for all $i\neq j$ and $\bigcup_{i}A_{i}=A$.
Definition 3.3 (Vector measure).
A vector measure on $\Theta$ with values in $\mathcal{Y}$ is a set function $\mu\colon\mathcal{B}(\Theta)\to\mathcal{Y}$ such that for all $A\in\mathcal{B}(\Theta)$ and all $\{A_{i}\}$ partitions of $A$,